OpenStudy (anonymous):

x/4+12=5 what is x? 3s-2s=60 what is s? -1/4m-3=5 what is m? 1/3y+10=2/3 what is y?

OpenStudy (anonymous):

x=-28 s=60 M=-2 Y=-28

OpenStudy (ash2326):

@yay1234 did you understand?

OpenStudy (ash2326):

@corey.tera please explain how did you solve it? Thanks

OpenStudy (anonymous):

yes an explanation would be apreciated i would like to know how to solve them so that i can do it myself if this comes up again

OpenStudy (anonymous):

x/4+12=5 -12 from both sides x/4=-7 then multiply both sides by 4 to get x alone so X= -28 3s-2s=60 Subtract like terms then you have 1s=60 So S=60 -1/4m-3=5 add three to both sides -1/4m=8 then multiply both sides by -1/4 M=-2 1/3y+10=2/3 Subtract 10 both sides 1/3y=-9.33333333333333 multiply both sides by 1/3 Y=-28

OpenStudy (anonymous):

Did that help a little more?

OpenStudy (ash2326):

Yeah Corey, you did a great work, Thanks:D

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Your Welcome

OpenStudy (mayliz):

x=28 s=60 m=-32 y=-28